### v1

# Upper critical field and pairing symmetry of Ising superconductors

Abstract
Motivated by the fact that the measured critical field $H_{c}$ in various transition metal dichalcogenide (TMD) superconductors is poorly understood, we reexamine its scaling behavior with temperature and spin-orbit coupling (SOC). By computing the spin-susceptibility in a multipocket system, we find that segments of the Fermi Surface (FS) at which the SOC has nodal points can have a contribution orders of magnitude larger than the remaining FS, hence setting the $H_{c}$, assuming the presence of a conventional singlet superconducting order parameter. Nodal lines of an Ising SOC in the Brillouin zone are imposed by symmetry, so they cause such nodal points whenever they intersect an FS pocket, which is indeed the case in monolayer NbSe2 and TaS2, but not in gated MoS2 and WS2. Our analysis reinterprets existing measurements, concluding that a dominant singlet-order parameter on pockets with SOC nodes is consistent with the $H_{c}(T)$ data for all monolayer Ising superconductors, in contrast to previous contradictory pairing assumptions. Finally, we show that the theory is also consistent with data on homobilayer TMDs.
Figure 1: a) FS of a monolayer Ising superconductor, of the same type as NbSe2. Two fully spin-polarized bands $\xi_{\pm}=\xi_{k}\pm\lambda(\boldsymbol{k})$ are split by the Ising SOC $\lambda(\boldsymbol{k})$. We model $\lambda^{\Gamma}(\theta)=\lambda^{\Gamma}_{0}\cos\left(3\theta\right)$ and $\lambda^{K}(\theta)=\lambda^{K}_{0}$, where $\lambda^{\Gamma}_{0}\approx\lambda^{K}_{0}/3$. b) The contribution to the difference in susceptibility between a uniform singlet superconducting state and the normal state. At the nodal points of the SOC $\lambda^{\Gamma}(\theta)=0$ the superconductor is a conventional singlet order. As the SOC is large everywhere else, $\delta\chi(T)$, and consequently the critical field $H_{c}$, is determined almost entirely by the 6 nodal points. c) & d) The normal and superconducting state susceptibilities $\chi_{N}(E),\chi_{S}(E)$ are shown on the $\Gamma$-pocket along radial lines, passing through a nodal point (c), or between nodal points (d). For spin-split bands in (d), the energy range where the combined susceptibility of the bands is non-zero extends to $\xi_{k_{\pm}}=\pm\lambda_{0}$ (shaded). In (c) the plots are individually scaled as max$|\chi_{N}(E)|\gg$max$|\chi_{S}(E)|$.
Introduction. For a given temperature, the upper critical field $H_{c}$ refers to the maximum magnetic field below which a material remains superconducting. For a singlet superconductor with spin-degenerate bands $H_{c}(T=0)$ is determined by the Pauli limit $H_{p}=1.76k_{\text{B}}T_{c}/\sqrt{2}\mu_{B}$ [1, 2]. However, it was long recognized that this Pauli limit can be exceeded in presence of the spin-orbit coupling (SOC)[3, 4, 5]. Many recent experiments on two-dimensional (2D) layered superconductors have measured a $H_{c}>H_{p}$. This is particularly noticeable in monolayer superconducting transition metal dichalcogenides (TMDs), also known as Ising superconductors, showing very high in-plane critical fields[6, 7, 8]. These materials are characterized by a large SOC originating from the broken inversion symmetry. Though remarkable, in these systems it remains difficult to extract significant information about the nature of the superconducting order parameter, such as being singlet or triplet [9, 10].
Additional puzzling issues arise in the 1H-MX2 compounds, M$=$ Nb, Ta, Mo, W, and X$=$ S, Se. The measured $H_{c}$ for monolayer NbSe2 and TaS2 indicates a square root scaling with the SOC while theories for a spin-split Fermi surface (FS) predict a linear scaling [4, 11, 12]. Moreover, if the bands are spin-split, the intervalley scattering required to fit the experimental data should be several orders of magnitude higher in NbSe2 than in WS2[8, 11], something for which there has been no physical explanation.
$\delta\chi\propto$ no SOC constant SOC nodal SOC
singlet $\Delta^{0}$ $\Delta^{2}$ $\Delta$
triplet, $\bm{h}\perp\bm{d}$ $\Delta^{2}$ $\Delta^{2}$ $\Delta^{2}$
$H_{c}/H_{p}\propto$ no SOC constant SOC nodal SOC
singlet 1 $\frac{\lambda}{\Delta}$ $\sqrt{\frac{\lambda}{\Delta}}$
triplet, $\bm{h}\perp\bm{d}$ $\frac{\epsilon}{\Delta}$ $\frac{\sqrt{\epsilon^{2}-\lambda^{2}}}{\Delta}$ $\frac{\sqrt{\epsilon^{2}-\lambda^{2}}}{\Delta}$
Table 1: The leading order term of the expansion of the susceptibility difference $\delta\chi$ in a small superconducting order parameter $\Delta$, for a magnetic field $\bm{h}\perp\bm{g}$ with $\bm{g}$ the spin-direction of the SOC (top). The resulting scaling with SOC strength $\lambda$ of the critical field, assuming a single pocket of the FS, with a given profile of SOC (absent; constant over the pocket; having nodal points). For a multi-pocket FS one pocket can have a $\delta\chi$ orders of magnitude larger than for the other pockets. Of note are pockets with a nodal SOC, where singlet pairing produces a unique scaling.
In this Letter, we calculate the critical field via the susceptibility taking into account the contribution of all pockets present at the Fermi surface. We find that the $\Gamma$-pocket, when present, almost entirely dominates the susceptibility and hence determines the critical field $H_{c}(T)$ in Ising superconductors. This is due to the fact that by symmetry the Ising SOC must have nodal points (at which there is no spin splitting) on the $\Gamma$-pocket [^1] Note that a Rashba SOC opens a small gap at these nodes, but our conclusion remains for a dominant Ising SOC, which is the case in TMDs. See [15] for details and references [55, 56] therein.. Despite the $H_{c}$ having been calculated decades ago for an isolated pocket with SOC nodes[14], its potential importance, nor the interplay with other pockets, has not been fully appreciated for superconducting TMDs[10]. For example, due to the position of its Fermi level, NbSe2 has the additional pocket around the $\Gamma$-point while WS2 does not. For a singlet pairing, the susceptibility at these nodal points is simply that of a conventional superconductor which therefore suppresses $H_{c}$. Our results on the scaling of the susceptibility for singlet and triplet order parameters are summarized in Table 1. We get values that are consistent with existing experimental data over all monolayer compounds. We find that the scaling of the critical field $H_{c}$ with SOC indicates that the effective pairing interaction between pockets is not negligible, and spin-singlet is the dominant order parameter in each compound in the family of monolayer Ising superconductors. Finally, we show that the theory matches the data on $2H-$type homobilayers, where the inversion symmetry is broken by the substrate.
Upper critical field from susceptibility. The critical field $H_{c}$ can be defined as the field at which the thermodynamic potentials $\Omega_{i}(T,H)$, for the normal state $(i=N)$ and for a proposed superconducting state $(i=S)$, are equal in value[2, 1]. Throughout, we will be assuming a temperature near the critical temperature $T_{c}$, so that $H_{c}$ is small enough to expand the potential in terms of the susceptibility. This standard procedure gives
$H_{c}(T)=\sqrt{\frac{\Omega_{0}(T)}{\delta\chi(T)}},$ (1)
where $\Omega_{0}(T)$ is the total condensation energy[15, 16], and $\delta\chi(T)=\chi_{N}(T)-\chi_{S}(T)$ is the difference in susceptibility between the two states. Near $T_{c}$, it is insightful to further expand the susceptibility difference in orders of the superconducting order parameter $\Delta$:
$\delta\chi(T)=\delta\chi_{0}(T)+\delta\chi_{1}(T)+\delta\chi_{2}(T)+\mathcal{O}\left(\Delta^{3}\right),$ (2)
where $\delta\chi_{m}=\mathcal{O}(\Delta^{m})$. The leading term then determines the value of the critical field, for example, if it is the zero-th order then $H_{c}(T)=\sqrt{\Omega_{0}(T)/\delta\chi_{0}(T)}\approx H_{p}$, while if it is the first order then $\sqrt{\Omega_{0}(T)/\delta\chi_{1}(T)}\gg H_{p}$. This connection is summarized for known cases in Table 1, which includes various scenarios for the SOC, which we discuss now.
SOC and multiple FS pockets in TMDs. We now consider the effective TMD transition metal bandstructure throughout the Brillouin zone (BZ), $\xi_{{\bm{k}},\zeta}=\xi_{\bm{k}}+\zeta\lambda(\bm{k})$, with $\zeta=\pm$ labeling the bands that are split by the SOC $\lambda(\bm{k})$. Note a key property: the Ising SOC is described by the scalar function $\lambda(\bm{k})$ that must be odd in $\bm{k}$, and hence must have nodal lines emanating from the $\Gamma$ point in the Brillouin zone.
Furthermore, for given bands and doping level, the FS can form multiple pockets, typically in TMDs around the $\Gamma$ and $K$ points in the BZ (see Fig. 1a). For each pocket, labeled $j=\Gamma,K\ldots$, we consider a parabolic band $\xi^{j}_{k}$ in the energy range ±$\pm\epsilon$ which captures the SOC splitting, namely, $\epsilon>\lambda^{j}_{0}$ [^2] The energy range must reach some physically relevant cut-off energy $\epsilon=\lambda_{0}^{j}+\epsilon_{c}$. In conventional superconductors with $\lambda_{0}=0$, the Debye frequency $\epsilon_{c}=\omega_{\text{D}}$ is chosen., where the SOC function at the pocket simplifies to $\lambda^{j}(\bm{k})=\lambda^{j}_{0}g^{j}(\theta)$, with $\theta$ the polar angle around the pocket and $g^{j}$ a dimensionless angular function.
The normal state susceptibility is a sum of contributions from each pocket, $\chi_{N}=\sum_{j}\chi_{N}^{j}$. We consider only in-plane magnetic fields $h$, so due to the two spin-split bands $\zeta=\pm$ at a pocket $j$, the susceptibility can further be usefully divided into two contributions of different physical origin. We have the intraband (Pauli) and the interband (van Vleck) contributions[18] $\chi^{j}_{N}=\chi_{N}^{j,\text{intra}}+\chi_{N}^{j,\text{inter}}$, with
$\chi_{N}^{j,\text{intra}}=$ $\sum_{\zeta,\bm{k}}\left(\left.\frac{\partial\xi^{j}_{\zeta,h}}{\partial h}\right|_{h=0}\right)^{2}\frac{\partial f(\xi^{j}_{\zeta})}{\partial\xi^{j}_{\zeta}}~{},$ (3)
$\chi_{N}^{j,\text{inter}}=$ $\sum_{\zeta,\bm{k}}\left.\frac{\partial^{2}\xi^{j}_{\zeta,h}}{\partial h^{2}}\right|_{h=0}f(\xi^{j}_{\zeta}),$ (4)
where $f$ the Fermi-Dirac distribution. The intraband term is strongly peaked at the FS, however due to the in-plane field any finite splitting $\lambda^{j}(\bm{k})\neq 0$ makes this term vanish. In contrast, the interband term has a priori contributions from all filled states in the pocket. However, we find cancellations everywhere except in the energy region containing the band splitting (see Fig. 1d). Physically, the finite splitting makes the susceptibility interband, originating from the canting of spins within the energy range of the splitting.
Susceptibility in the superconductor. Throughout we will consider only pairing between $\bm{k}$ and $-\bm{k}$ electrons, hence we can define the pairing functions $\Delta^{j}$, for example, for the $j=\Gamma$ pocket, and $j=K$ for pairing between $K$ and $-K$ pockets. We will assume that the temperature dependence of the pairing functions $\Delta^{j}$ follows the mean-field solution for $h=0$ with $\Delta^{j}(T=0)=\Delta^{j}_{0}=1.76k_{\text{B}}T^{j}_{c}$. We will discuss in detail the relationships between different $\Delta^{j}$ in the next section.
Focusing on one pairing $\Delta^{j}$, the condensation energy is actually independent of the SOC $\lambda^{j}_{0}$[15]. Thus, the condensation energy $\Omega^{j}_{0}(T)$ is simply proportional to the density of states of the $j$ pocket and to $(\Delta^{j}_{0})^{2}$ (see [15]).
The susceptibility in the superconducting state also has two contributions for each fixed $j$, $\chi^{j}_{S}=\chi_{S}^{j,\text{intra}}+\chi_{S}^{j,\text{inter}}$, defined for the BdG bands $E^{j}_{\zeta}$ as[19]:
$\chi_{S}^{j,\text{intra}}=\sum_{\zeta,\bm{k}}\left(\left.\frac{\partial E^{j}_{\zeta,h}}{\partial h}\right|_{h=0}\right)^{2}\frac{\partial f(E^{j}_{\zeta})}{\partial E^{j}_{\zeta}},$ (5)
$\chi_{S}^{j,\text{inter}}=$ $\sum_{\zeta,\bm{k}}\left(\left.\frac{\partial^{2}E^{j}_{\zeta,h}}{\partial h^{2}}\right|_{h=0}f(E^{j}_{\zeta})\right.$ (6)
$\left.+\frac{1}{2}\left(\left.\frac{\partial^{2}\xi^{j}_{\zeta,h}}{\partial h^{2}}\right|_{h=0}-\left.\frac{\partial^{2}E^{j}_{\zeta,h}}{\partial h^{2}}\right|_{h=0}\right)\right).$
Let us consider three revealing examples. First, in a conventional superconductor, with spin-singlet s-wave pairing and spin-degenerate bands (so $\lambda_{0}^{j}\equiv 0$), the $\chi^{j}_{N}(T=0)=\chi_{P}$ while $\chi^{j}_{S}(T=0)=0$. Second, if the same bands have a triplet pairing function characterized by the vector $\boldsymbol{d}\perp\boldsymbol{h}$, then $\chi^{j}_{S}(0)=\chi_{P}+\mathcal{O}\left(\Delta^{2}\right)$[4, 18]. Third, for either of the two previous pairing scenarios, but with spin-split bands [^3] Note that the spin-direction of Ising SOC is perpendicular to the in-plane magnetic field, otherwise the following expressions should be modified., $\lambda^{j}(\bm{k})\neq 0$, the $\chi^{j}_{S}(0)=\chi_{P}+\mathcal{O}\left(\Delta^{2}\right)$[18, 4]. As summarized in Table 1, the leading order of $\delta\chi^{j}$ (Eq. (2)) is highest (quadratic in $\Delta^{j}$) for triplets, or for fully split bands, giving the highest $H_{c}$. This leading order can be lowered if the splitting vanishes at some points on the pocket (i.e., there are SOC nodes $g^{j}(\theta_{node})=0$ for some $\theta_{node}$, see Fig. 1b), or even further by completely removing the splitting $\lambda_{0}^{j}=0$. Therefore, for a fixed pocket $j$, we find a rule-of-thumb: The upper critical field is lowest for the least exceptional scenario (no band splitting, singlet pairing).
Figure 2: Calculated upper critical field compared to experimental data for monolayer Ising superconductors with a $\Gamma$-pocket (NbSe2[6] & TaS2[8]) and without (gated MoS2[21] & WS2[22]). The $\Gamma$-pocket introduces a $\sqrt{\lambda_{0}/\Delta_{0}}$-scaling of $H_{c}$ compared to the otherwise linear $\lambda_{0}/\Delta_{0}$-scaling in the other compounds. The ratio $\lambda_{0}/\Delta_{0}$ is taken at the $K$-pockets.
Multipocket singlet superconductivity in TMDs. Our key observation now is that a $\Gamma$ pocket in Ising SOC TMDs must have SOC nodes, due to the SOC nodal lines. In contrast, the $K$ pockets do not (see Fig. 1a). Hence, for singlet pairings $\Delta^{j}$, reference systems with isolated pockets would have very different behaviors of $H^{j}_{c}$ (see [15] for definitions), raising the question of the behavior of the (unique) $H_{c}$ of the system.
We hence solve the self-consistent gap equations (see [15, 23]) for singlet s-wave pairings, with an additional simplifying assumption based on experimental data (at least in NbSe2 [24, 25, 26]): the pairings $\Delta^{\Gamma}$ and $\Delta^{K}$ are equal in size for all temperatures near $T_{c}$. (We also consider adding a triplet component, which turns out to be small near $T_{c}$, see [15].) We find that these assumptions self-consistently imply that inter-pocket pairing interaction is as strong as the intra-pocket one (see [15] for all details), implying that the superconductivity is strongly coupling the pockets. With such coupling, we get that very approximately:
$\frac{1}{H_{c}^{2}}\approx\frac{1}{3}\left(\frac{1}{(H^{\Gamma}_{c})^{2}}+2\frac{1}{(H^{K}_{c})^{2}}\right).$ (7)
If reference systems with isolated pockets $j$ would have critical fields $H_{c}^{j}$ differing by a factor of at least 2, the self-consistent $H_{c}$ of the three-pocket system would be set by the lower of the two $H_{c}^{j}$. We hence extend the previous rule-of-thumb: The self-consistent $H_{c}$ of a singlet Ising SOC TMD superconductor is set by the pocket $j$ with the lowest $H_{c}^{j}$, i.e., by the least exceptional part of the FS.
Monolayer. With the above insights, let us present in detail the model of monolayer NbSe2 with three pockets (two $K$-pockets and one $\Gamma$-pocket of equal size).
[…]
$\delta\chi^{K}(T)\approx\frac{\chi_{P}}{2}\frac{\Delta^{2}}{\lambda_{0}^{2}}\ln\left[\frac{\lambda_{0}^{2}}{\Delta^{2}}\right].$ (8)
Thus, as known previously[11, 12, 10, 21, 28, 22], if only $K$-pockets are considered the critical field scales linearly with the SOC $\frac{H^{K}_{c}(T)}{H_{p}}\propto\frac{\lambda^{K}_{0}}{\Delta^{K}_{0}}$. In fact, for the $K$-pockets both a singlet and a triplet pairing $\Delta^{K}$ give $\delta\chi^{K}(T)=\delta\chi_{2}^{K}(T)$.[4]
For the $\Gamma$-pocket we model $\lambda^{\Gamma}(\bm{k})=\lambda^{\Gamma}_{0}\cos\left(3\theta\right)$[29, 30], where $\theta$ is the polar angle around the $\Gamma$ point. For segment of the pocket near a SOC node (always assuming temperature near $T_{c}$) we determine that the difference in susceptibility is well-described by a Lorentzian $L(\theta,\beta)$ of width $\beta$ centered on the node $\theta=\theta_{0}$ (see [15] for details, and Fig. 1b):
$\delta\chi^{\Gamma}(\theta,T)\approx\left(\chi_{P}-\chi_{\text{sg}}(T)\right)\frac{\pi\Delta^{\Gamma}_{0}}{\sqrt{3}\lambda^{\Gamma}_{0}}L\left(\theta,\frac{\Delta^{\Gamma}_{0}}{\sqrt{3}\lambda^{\Gamma}_{0}}\right),$ (9)
with $\chi_{\text{sg}}(T)=\chi_{P}Y(T)$ is given by the Yoshida function $Y(T)$ characteristic of a singlet[18]. This contribution from near a node, $\theta\approx\theta_{0}$, is dominant on the $\Gamma$ pocket, and we find that it closely corresponds to the numerical result for the entire pocket when $T>0.8T_{c}$ (see [15] for a detailed comparison). Using this expression to integrate over the entire pocket, we get
$\delta\chi^{\Gamma}(T)=\left(\chi_{P}-\chi_{\text{sg}}(T)\right)\frac{\pi\Delta_{0}}{\sqrt{3}\lambda_{0}}+\mathcal{O}(\Delta_{0}^{2}),$ (10)
resulting in a critical field which now scales as a square root of the SOC:
$H^{\Gamma}_{c}(T)\approx\sqrt{\frac{\lambda^{\Gamma}_{0}}{\Delta^{\Gamma}_{0}}}\sqrt{\frac{\Omega^{\Gamma}_{0}(T)}{\left(\chi_{P}-\chi_{\text{sg}}(T)\right)\frac{\pi}{\sqrt{3}}}}\propto\sqrt{\frac{\lambda^{\Gamma}_{0}}{\Delta^{\Gamma}_{0}}}.$ (11)
Hence, for the $\Gamma$-pocket the assumption of a singlet is crucial: The SOC nodes on the $\Gamma$ pocket result in $\delta\chi^{\Gamma}(T)\gg\delta\chi^{K}(T)$ for singlet pairing and an in-plane field.
For the full 3-pocket model of NbSe2, $\Omega_{0}(T)$ and $\delta\chi(T)$ are determined by the ratios $\lambda_{0}^{K}/\Delta_{0}^{K}$, $\lambda_{0}^{\Gamma}/\Delta_{0}^{\Gamma}$, and $\Delta_{0}^{K}/\Delta_{0}^{\Gamma}$. As the superconductivity in NbSe2 is relatively uniform for few layers[24, 25, 26], we have assumed that $\Delta_{0}^{K}\approx\Delta_{0}^{\Gamma}$. Due to the large SOC in these compounds, $\delta\chi^{\Gamma}(T)\gg\delta\chi^{K}(T)$ and therefore $\delta\chi(T)\approx\delta\chi^{\Gamma}(T)$, in accord to the general discussion above. In Fig. 2 measurements and calculations for NbSe2 and TaS2, with $\frac{\lambda_{0}^{\Gamma}}{\Delta_{0}^{\Gamma}}\approx\frac{\lambda_{0}}{3\Delta_{0}}$[31, 32], are compared to TMDs where only $K$-pockets have been considered [^5] See [15] for a discussion of the additional $Q$-pockets present in gated MoS2[57, 58, 59, 60, 61]). This compound should also be considered a multipocket compound. However, both types of pockets are split by a constant SOC and the susceptibility difference is dominated by the pockets with the lowest value..
The result to highlight here is that the $\sqrt{\lambda_{0}/\Delta_{0}}$-scaling only appears if the superconducting order parameter has a finite singlet gap at the SOC nodes. For a triplet order a $\lambda_{0}/\Delta_{0}$-scaling is regained (see [15] for the critical field obtained for different pairing symmetries on a $\Gamma$-pocket). The fact that the experimental data for monolayer NbSe2 and TaS2 are related by $\sqrt{\lambda_{0}/\Delta_{0}}$-scaling (Fig. 2) is therefore a strong indicator that the superconducting order is mainly a singlet at the SOC nodes.
Figure 3: Mono- and bi-layer NbSe2[6] and TaS2[8, 7]. For bilayers a fit for the interlayer bias potential $\delta\mu$ is shown. For multilayers a lower bound of $H_{c}(T)$ is Eq. (12), the suppression from hopping between the layers at $\delta\mu=0$. The upper bound considers the layers as uncoupled monolayers at the given $T_{c}$ ($\delta\mu\rightarrow\infty$).
Multilayer. In multilayer TMDs, individual layers break inversion symmetry, yet in the 2H-stacking the overall inversion is preserved, and the sign of the SOC alternates between layers. Layers are coupled by an interlayer hopping $t_{\perp}$. This coupling is stronger for the $\Gamma$-pocket[34] due to the increased presence of $d_{z^{2}}$-orbitals[35, 36]. Motivated by this, all pockets should be included also in the multilayer model.
For intralayer singlet superconductivity the susceptibility can never reach beyond the value of the normal state van Vleck-susceptibility[37, 9, 38], as the remaining Pauli susceptibility originates from the Fermi surface. In a bilayer 2H-TMD the SOC nodes disappear and instead two sets of degenerate bands are separated by a distance proportional to $t_{\perp}$. In our model we take the same $t_{\perp}$ for all pockets. The now leading term $\delta\chi^{j}_{0}$ is determined by the ratio $t_{\perp}/\lambda_{0}^{j}$. For $K$- and $\Gamma$-pockets respectively, we find
$\delta\chi^{K,\text{2L}}(T)\approx\frac{t_{\perp}^{2}}{t_{\perp}^{2}+\left(\lambda_{0}^{K}\right)^{2}}(\chi_{P}-\chi_{\text{sg}}(T)),$ (12)
$\delta\chi^{\Gamma,\text{2L}}(T)\approx\frac{t_{\perp}}{\sqrt{t_{\perp}^{2}+\left(\lambda_{0}^{\Gamma}\right)^{2}}}(\chi_{P}-\chi_{\text{sg}}(T)).$ (13)
For any number of layers $N$ the leading term is $\delta\chi^{j}_{0}$, for up to $N=5$ layers[15].
Another non-negligible factor in the multilayer case is a small inversion symmetry breaking coming from a substrate. In the bilayer, we hence include an interlayer bias potential[39], $\delta\mu$. At the FS this splits the spin-polarized bands, and the SOC nodes reemerge. The inversion symmetry broken bilayer can therefore be modeled as two monolayer bands with an effective SOC $\tilde{\lambda}_{0}$. From a linearization of the bands close to the SOC nodes,
$\tilde{\lambda}^{j}_{0}=\frac{\delta\mu}{\sqrt{\delta\mu^{2}+t_{\perp}^{2}}}\lambda^{j}_{0}.$ (14)
If $\delta\mu$ is sufficiently large, the critical field thus scales as $H_{c}^{2L}/H_{p}^{2L}\propto\sqrt{\frac{\tilde{\lambda}^{\Gamma}_{0}}{\Delta_{0}^{\Gamma,\text{2L}}}}$. More generally, compared to the monolayer, the bilayer $H_{c}$ depends on an increased number of parameters: $\lambda^{j}_{0},\Delta_{0}^{j,\text{2L}},t_{\perp},\delta\mu$ for each pocket $j$. In results of Fig. 3 the value $\delta\mu$ has been fitted to experimental data, while $\lambda^{j}_{0},t_{\perp}$ are given by Ref. [8], and $\Delta_{0}^{j,\text{2L}}$ by $T_{c}$ for the fitted dataset. For both NbSe2 and TaS2 the find $\delta\mu<10$meV, which is reasonable [40, 41, 39, 42, 43].
To deal with multiple model parameters, we consider an upper and lower bound for the $N$-layer $H_{c}$. The lower bound is given by Eq. (12), where the sample is considered fully isolated from its environment. The upper bound is the limit where each layer is treated as an isolated monolayer. For layers $N>2$, the bounds can be calculated (see [15] for details), while additional parameters are required to model the effects of a substrate.
In contrast to the monolayer case the multilayer $H_{c}$ does not carry a strong indication about the pairing symmetry at specific points in momentum space. However, for a fully triplet order the interlayer hopping would not suppress $H_{c}$[37, 9].
Conclusions. We have demonstrated the necessity of including all bands when calculating the upper critical field of a superconductor. The critical field can be limited by the sections of the FS where $\bm{g}$ or $\bm{d}$ is no longer perpendicular to the field. We have illustrated this principle by comparing $H_{c}$ for Ising SOC superconductors either having or not pockets where the SOC has nodal points. We have shown a huge difference in $H_{c}$ arising for different pairing symmetries at these few SOC-node points. Therefore, the behavior of the upper critical field could be used to indicate which pairing symmetries at specific points in momentum space are favored.
For monolayer 1H-NbSe2 and 1H-TaS2, our analysis shows that a dominant triplet order parameter is inconsistent with the experimental data. This rules out any purely $f$-wave[44, 45] or nematic order[46, 47, 48]. Furthermore, even though our calculation only includes a uniform singlet order, other proposed orders for NbSe2 are consistent with our results. This includes a nodal singlet, a topologically non-trivial singlet order such as a $d+id$-wave, or the recently proposed nodal topological superconductor[29, 44, 49, 50, 51]. A mixed parity order, $s+f$-wave[52, 53] or a $d_{y}$-triplet component appearing under a field[54, 10, 24], would also only affect the subleading order of $\delta\chi(T)$ and thus only slightly change the value of $H_{c}$. It should also be reiterated that the signature from the $K$-pockets are similarly almost entirely hidden, as their potential influence on $H_{c}$ is of the same size as disorder effects[11, 10].
Other works have also found the $\Gamma$-pocket essential to accurately describe a pairing mechanism dependent on spin-fluctuations[48], reinforcing the conclusion that the nodal points can play a vital role.
Note that a Rashba SOC $\lambda_{R}$, e.g., due to a substrate, lifts the Ising SOC nodes. We show that the correction to the susceptibility difference is only of order $(\lambda_{R}/\lambda_{0})^{2}$, and hence the influence on $H_{c}$ is negligible (see [15]).
In this work, we focused on Ising SOC superconductors as characteristic inversion-symmetry-breaking materials with only a few points in the FS without a SOC splitting. However, the idea is general as it relies on symmetries and shows that the critical field could similarly be used to extract information on the nature of the order parameter in hetero- and/or Moiré structures.
Acknowledgment. We thank fruitful discussions with M. Calandra, C. Quay and I. Paul. This work was supported by the French Agence Nationale de la Recherche (ANR), under grant number ANR-22-CE30-0037.